Black pencil cost#15 each and colored pencil cost #21 each if24 mixed pencil costcost #402 how many of them were black
Answer by JBarnum(2146)

b=black
c=colored
b+c=24
15b+21c=402
Use substitution
b=24-c
15(24-c)+21c=402
360-15c+21c=402
360+6c=402
6c=42
c=7
b=24-7
b=17
17 black pencils